OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

hoWWWWW

Vocaloid (vocaloid):

first find the magnitude of the vector u

Vocaloid (vocaloid):



Vocaloid (vocaloid):

so, what is the magnitude of vector u?

OpenStudy (amistre64):

if an object is 10 units long, then 1/10th of it is 1 unit long if an object is 3 units long, then 1/3rd of it is 1 unit long if an object is 1/5 units long, then 5/1ths of it is 1 unit long if an object is L units long, then 1/Lth of it is 1 unit long

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

If, \(\Large\color{black}{ \displaystyle \vec{v}= \langle a,b\rangle }\) Then, the unit vector in direction of v is, \(\Large\color{black}{ \displaystyle \langle \frac{a}{\left|\vec{v}\right|},{~}\frac{b}{\left|\vec{v}\right|}\rangle }\) (divide each component by the vector's magnitude.

OpenStudy (anonymous):

wait so

OpenStudy (anonymous):

-4/5,-3/5 ????

OpenStudy (solomonzelman):

yes, that is correct.